How do I write the equation of the line that has a y-intercept of (0,-6) and is perpendicular to the line 3x - 6y = 8.
Answer by nerdybill(7384)

How do I write the equation of the line that has a y-intercept of (0,-6) and is perpendicular to the line 3x - 6y = 8.
.
Remember the "slope-intercept" form:
y = mx + b (equation of a line)
.
From:"the line that has a y-intercept of (0,-6)" we now know:
b = -6
.
To find 'm', begin by finding slope of:
3x - 6y = 8
-6y = -3x+8
6y = 3x-8
y = (3/6)x-(8/6)
y = (1/2)x-(4/3)
a line that is perpendicular to the above line MUST have a slope that is the negative reciprocal:
(1/2)m = -1
m = -2
.
Plug 'm' and 'b' back into:
y = mx + b
to get your answer:
y = -2x + (-6)
y = -2x - 6